OpenStudy (anonymous):

f(x) = cos(x - 4) - 8 find the inverse function of the function f

OpenStudy (anonymous):

y= cos(x-4) -8 cos(x-4) = y+8 x-4 =cos^{-1}(y+8) x=4+ cos^{-1}(y+8) so the inverse is 4+ cos^{-1}(y+8)

myininaya (myininaya):

y=cos(x-4)-8 (you should define a domain for this function that will make the function 1 to 1) y+8=cos(x-4) cos^{-1}(y+8)=x-4 cos^{-1}(y+8)+4=x f^{-1}(x)=cos^{-1}(x+8)+4

OpenStudy (anonymous):

sin(cos-1 1/2 - sin-1 squareroot 3 over 2) --> how do you find the exact value of this??

myininaya (myininaya):

cos^{-1}(1/2)=a cosa=1/2 so a=pi/3 sin^{-1}(sqrt{3}/2)=b sinb=sqrt{3}/2 so b=pi/3 sin(pi/3-pi/3)=sin(0)=0

OpenStudy (anonymous):

dang thanks!!

OpenStudy (anonymous):

cos(tan-1 5/12 - cos-1 4/5) its multiple choice A) 7/13 B)63/65 c)52/65 d)13/24

OpenStudy (anonymous):

thats simple

myininaya (myininaya):

B

OpenStudy (anonymous):

let a be the first thing, b be the second and expand

myininaya (myininaya):

also you can just put it in you calculator and then look at the fractions to see which one it is

myininaya (myininaya):

thats the cheating way

OpenStudy (anonymous):

sin[pi/4 + x] = squareroot2(cos x + sin x)

OpenStudy (anonymous):

I'm suppose to establish the identity

OpenStudy (anonymous):

you can express tan inverse in terms of cos inverse as tan^(-1) (a/b)= cos^(-1) (b/sqrt(a^2+b^2)) then use cos^(-1) a + cos^(-1) b formula to find the answer

OpenStudy (anonymous):

thanks shankvee!!!!! for all your help!